# Sloth: scaling laws for LLM skills to predict multi-benchmark performance across families

### v3

@@ Abstract
Scaling laws for large language models (LLMs) predict model performance based on parameters like size and training data. However, differences in training configurations and data processing across model families lead to significant variations in benchmark performance, making it difficult for a single scaling law to generalize across all LLMs. On the other hand, training family-specific scaling laws requires training models of varying sizes for every family. In this work, we propose Skills Scaling Laws (SSLaws, pronounced as Sloth), a novel scaling law that leverages publicly available benchmark data and assumes LLM performance is driven by low-dimensional latent skills, such as reasoning and instruction following. These latent skills are influenced by computational resources like model size and training tokens but with varying efficiencies across model families. Sloth exploits correlations across benchmarks to provide more accurate and interpretable predictions while alleviating the need to train multiple LLMs per family. We present both theoretical results on parameter identification and empirical evaluations on 12 prominent benchmarks, from Open LLM Leaderboard v1/v2, demonstrating that Sloth predicts LLM performance efficiently and offers insights into scaling behaviors for downstream tasks such as coding and emotional intelligence applications. Our code is available on https://github.com/felipemaiapolo/sloth.
@@ 1 Introduction
In this work, we propose a new class of scaling laws called Sloth to solve this dilemma. These scaling laws are fitted using publicly available data (e.g., from LLM leaderboards) across multiple benchmarks, leveraging information shared among benchmarks and model families to improve prediction power and interpretability through parameter efficiency, i.e., fewer parameters without hurting performance. Specifically, we utilize the correlations in benchmark scores to make the scaling law simpler in terms of parameter count without harming prediction power by assuming that LLM performance is driven by a set of low-dimensional latent skills, such as reasoning and instruction following, which can be easily interpreted. Furthermore, we hypothesize that these latent skills are similarly influenced by computational resources, such as model size and training tokens, across different LLM families, with the key distinction being each family’s efficiency in converting compute into skill levels. In summary, our main contributions are
@@
  • Providing a theoretical result regarding the identification of Sloth’s parameters and empirically demonstrating that our scaling laws can (i) accurately predict the performance of large models in 12 prominent LLM benchmarks and (ii) provide interpretable insights into LLM scaling behavior.
  • Demonstrating how predicted latent skills can be used to predict model performance in complex downstream tasks such as coding and emotional intelligence applications.
@@ 1.1 Related work
Scaling laws for deep neural networks: In recent years, researchers have studied scaling laws from different angles. Rosenfeld et al. (2019) provides experimental scaling laws that predict model loss as a function of training set size, model width, and model depth. Likewise, Kaplan et al. (2020) establishes scaling laws that primarily measure loss (perplexity) and not accuracy on downstream tasks or benchmarks. Motivated by the presence of hard limits on the size of trainable data sets but a hypothetical unlimited ability to scale models, the authors of Muennighoff et al. (2023) establish scaling laws in constrained data settings. They find that perhaps unsurprisingly, increasing computing provides diminishing returns if data does not scale. Gadre et al. (2024) addresses the gap between the assumptions in scaling laws and how training is performed in practice; in particular, they construct scaling laws that both perform well in the over-training regime and predict performance on downstream tasks. In a similar but distinct direction, some works try not only to estimate scaling laws but also respond to the following strategic question: “Given a fixed FLOPs budget, how should one trade-off model size and the number of training tokens?” For example, Hoffmann et al. (2022) provides a partial answer, introducing the celebrated family of Chinchilla scaling laws and finding that training tokens and parameter size should roughly scale together. This contrasts with the older work of Kaplan et al. (2020) that provides a series of power laws that imply that simply increasing parameter count will provide good returns. Each of these referenced works trains models with a particular pretraining setting (e.g., architecture) at various sizes and ultimately seeks to predict test loss. Our focus is distinct, we fit scaling laws on existing benchmark data of multiple model families and predict LLM benchmark performance with minimal amount of data on the new family being predicted. The closest related works are Owen (2024); Ruan et al. (2024); Gadre et al. (2024); we will provide a detailed comparison with their work throughout the paper.
@@
LLMs latent skills: Given that the performance of large language models (LLMs) in different and diverse benchmarks is correlated, it makes sense to think that those models have some low-dimensional latent skills that are reflected in downstream tasks. In this direction, Ilić (2023) extracts a general intelligence factor (“g-factor”) for LLMs using the Open LLM Leaderboard (Beeching et al., 2023) and GLUE (Wang et al., 2018) using factor analysis. They also verify that this “g-factor” positively correlates with model size. In a similar direction Burnell et al. (2023) uses HELM (Liang et al., 2022) data to reveal that LLM intelligence may be constituted by three distinct, yet correlated factors. They also verify a positive correlation between model size and these latent skills, yet they do not propose a formal scaling law. In their study, the authors do not account for the training set size or model family information, leading to a poor fit of the regression model; this leaves good extrapolation as an open problem we address. In Kipnis et al. (2024), a unidimensional item response theory model is applied to each one of the 6 (filtered) benchmarks of the Open LLM Leaderboard. A factor analysis on the skill parameters shows that the main factor (carrying  80% of the data variability) is highly correlated with the “grand” (average) score of LLMs. In a related but different direction, Maia Polo et al. (2024a; b) show that inferring low-dimensional latent skills of LLMs can make model evaluation much more efficient, saving up to 140x in computing power. In this work, we explicitly model LLM skills as a function of computing resources, which enables the creation of accurate and interpretable scaling laws for benchmark performances.
@@ 3.1 Model architecture
We present a novel scaling law called Sloth, which introduces several modifications to (2.2). The key innovation of Sloth lies in its explicit modeling of the correlation structure between benchmarks, resulting in improved predictive accuracy and interpretability. Moreover, Sloth proposes that (i) LLM capabilities should scale with computing resources similarly across families up to an efficiency factor, (ii) benchmark performance can depend on $s$ and $t$ not only through the total number of FLOPs, and (iii) that the function $\sigma$ can also be learned in cases in which predictive performance is important. We detail each one of these points.
Inspired by the latent skills (e.g., reasoning, language modeling, instruction following) inferred from benchmark data in Burnell et al. (2023); Ilić (2023); Ruan et al. (2024); Gor et al. ; Kipnis et al. (2024); Maia Polo et al. (2024a; b), we propose creating a scaling law for LLMs skills by leveraging the correlation structure of the benchmarks; for example, we model how the construct “reasoning” scales with compute instead of modeling benchmarks scores directly. The two major advantages of this approach are better performance prediction since we have fewer parameters to fit (reducing overfitting) and extra interpretability/insights. Concretely, we model $\eta_{ij}(s,t)$’s simultaneously for benchmarks $j\in\mathcal{J}$ as each being a linear combination of the same low-dimensional latent skills $\theta_{i}(s,t)\in{\mathbf{R}}^{d}$ plus a bias term $b\in{\mathbf{R}}^{J}$, where $d\ll J=|\mathcal{J}|$. Denote $\eta_{i}(s,t)\in{\mathbf{R}}^{J}$ as the vector of $\{\eta_{ij}(s,t)\}_{j\in\mathcal{J}}$. Mathematically, we have
@@
One can see that $\Lambda\in{\mathbf{R}}^{J\times d}$ encodes the correlation structure between the benchmarks; in particular, it tells us which benchmark measures overlapping (or distinct) skills. Interestingly, our model has a strong connection with factor analysis (FA) models, which we elaborate on in detail in Appendix C. In FA, the matrix $\Lambda$ is known as factor loadings while $\theta_{i}(s,t)$ are known as factors.
@@
Note that (i) the intercept parameter $\alpha_{ik}$ is indeed family-dependent while each skill slope is shared across families and (ii) $\theta_{i}$ can depend on $s$ and $t$ not only through $c(s,t)$. In economic terms, the intercept term $\alpha_{ik}$ can be interpreted as an efficiency measure of the family $i$ in converting compute to performance for skill $k$ and, in practice, will absorb all hidden factors specific to family $i$ such as data quality, post-training factors, etc.. We note that the interaction term in ($\log(s)\log(t)$) accounts for the fact that the impact of $\log(s)$ and $\log(t)$ on skills might depend on each other; in Appendix D, we show some evidence that this is indeed the case. Additionally to the changes in $\eta_{ij}$, we propose making the activation function $\sigma$ trainable and specific to each benchmark $j$ if needed. To that end, we adopt a semi-parametric single-index model using neural networks (Bietti et al., 2022). To make the results more behaved if (out-of-support) generalization is needed, we assume $\sigma_{j}:{\mathbf{R}}\to[0,1]$ is given by a monotonic (increasing) neural network, which can be achieved by constraining its weights to be non-negative (Sill, 1997) and its last activation function to be sigmoid. We note, however, that one can always forgo training of the link function and instead assume a sigmoid structure as this simplifies model fitting and may make the model more interpretable.
3.2 Identifiability of model parameters
Assumption 3.1 (Identifiability constraints).
Assume that
$\frac{1}{n}\sum_{i=1}^{n}\theta^{(i)}=0,\quad\frac{1}{n}\sum_{i=1}^{n}{\theta^{(i)}}^{\top}\theta^{(i)}=\Psi,\quad\text{rank}(\Lambda)=d,\quad\text{and}\quad\text{rank}(X)=p,$
where $\theta^{(i)}$ denotes the $i$-th row of $XB$ and $\Psi$ is a positive definite matrix.
One possible choice for the covariance matrix is $\Psi=I_{d}$ (Chen et al., 2019), which assumes uncorrelated skills. One implicit implication of Assumption 3.1 is that $n\geq p\geq d$ must be satisfied, otherwise the covariance matrix cannot be full rank. This condition is satisfied in our experiments. Under Assumption 3.1, we show the identifiability of the model parameters up to a transformation of $\Lambda$ tied to a transformation of $B$, which leaves the outputs of the model unchanged. This means that we can potentially approximate the true values for $\Lambda$ and $B$ up to a transformation, which is usually the norm within the class of exploratory factor analysis models.
Theorem 3.2.
then, under the Assumption 3.1, we have $\tilde{b}=b$, $\tilde{\Lambda}=\Lambda M$, and $\tilde{B}=B(M^{\top})^{-1}$ for an invertible matrix $M\in{\mathbf{R}}^{d\times d}$. In particular, $M$ is orthogonal if $\Psi=I_{d}$, i.e., $M^{\top}M=MM^{\top}=I_{d}$.
We place the proof of Theorem 3.2 in Appendix B. From our proof, we can see that the matrix $M$ is dependent on the specification of $\Psi$.
3.3 Model fitting
[…]
@@ 3.4 Interpretability and practical considerations post model fitting
In practical situations, it is hard to fix the covariance matrix of skills to something meaningful before fitting the model, as suggested in Section 3.1. To make the model interpretable, we mirror a standard approach used in factor analysis, e.g., in Chen et al. (2019)’s applications. First, we fit Sloth without any constraints on the covariance of skills obtaining the estimates $(\hat{\Lambda},\hat{b},\hat{B})$. Second, we find the matrix $A\in{\mathbf{R}}^{d\times d}$ such that the skills $X\hat{B}A$ have covariance identity, update $\hat{B}\leftarrow\hat{B}A$, and update $\hat{\Lambda}\leftarrow\hat{\Lambda}(A^{\top})^{-1}$ so the model outputs remains unchanged, because $\hat{\Lambda}(X\hat{B})^{\top}=\hat{\Lambda}(A^{\top})^{-1}(X\hat{B}A)^{\top}$. Third, we find a matrix $M\in{\mathbf{R}}^{d\times d}$ such that $\hat{\Lambda}M$ is easily interpretable (e.g., it is a sparse matrix); there are different methods to find $M$ and, in this paper, we use the Geomin (Yates, 1987; Chen et al., 2019) oblique rotation method to find a suitable $M$ using the Python package FactorAnalyzer (Biggs, 2019). We then update $\hat{\Lambda}\leftarrow\hat{\Lambda}M$ and, to make the model invariant, we also update $\hat{B}\leftarrow\hat{B}(M^{\top})^{-1}$; the resulting skills are still guaranteed to have unitary standard deviations, so their covariance equals their correlation. Finally, we standardize the columns of the skills $X\hat{B}$ to have zero mean, while keeping the correlation structure unchanged. This last step implies that $\hat{b}$ must be translated to make the model invariant.
@@ 4 Sloth in practice
In this section, we present some experimental results that provide evidence of the usefulness of Sloth. We perform experiments on a set of twelve benchmarks and state-of-the-art LLM families, including LLaMa 3 (Dubey et al., 2024), Qwen 2 (Yang et al., 2024), and Yi 1.5 (Young et al., 2024). We explore the following applications: (i) benchmark performance prediction for larger models from a specific LLM family, (ii) interpretability of the scaling of skills (can help practitioners allocate resources based on the skills of interest), and (iii) complex downstream tasks performance prediction.
@@ 4.1 Data
We expand the dataset made available by Ruan et al. (2024), including more models from the HuggingFace Open LLM leaderboard v1 (Beeching et al., 2023) and v2 (Fourrier et al., 2024). In our extended dataset, we have a total of 30 families [^1] If we consider that instruct and base models are from different families, we end up with 53 families., which 28 are on v1 of the Open LLM Leaderboard and 17 families measured on v2 of the Open LLM Leaderboard. Furthermore, there are 15 families at the intersection of the two versions. Furthermore, we collect data and present results on the performance of a variety of instruction-tuned versions of the base models we consider. As far as we are aware, our dataset is the most comprehensive among prior works on benchmark data scaling laws. Please check Appendix F for details on the included models.

### v4

@@ Abstract
Scaling laws for large language models (LLMs) predict model performance based on parameters like size and training data. However, differences in training configurations and data processing across model families lead to significant variations in benchmark performance, making it difficult for a single scaling law to generalize across all LLMs. On the other hand, training family-specific scaling laws requires training models of varying sizes for every family. In this work, we propose Skills Scaling Laws (SSLaws, pronounced as Sloth), a novel scaling law that leverages publicly available benchmark data and assumes LLM performance is driven by low-dimensional latent skills, such as reasoning and instruction following. These latent skills are influenced by computational resources like model size and training tokens but with varying efficiencies across model families. Sloth exploits correlations across benchmarks to provide more accurate and interpretable predictions while alleviating the need to train multiple LLMs per family. We present both theoretical results on parameter identification and empirical evaluations on 12 prominent benchmarks, from Open LLM Leaderboard v1/v2, demonstrating that Sloth predicts LLM performance efficiently and offers insights into scaling behaviors for complex downstream tasks and increased test-time compute. Our code is available on https://github.com/felipemaiapolo/sloth.
Machine Learning, ICML
@@ 1 Introduction
In this work, we propose a new class of scaling laws called Sloth to solve this dilemma. These scaling laws are fitted using publicly available data (e.g., from LLM leaderboards) across multiple benchmarks, leveraging information shared among benchmarks and model families to improve prediction power and interpretability through parameter efficiency, i.e., fewer parameters without hurting performance. Specifically, we utilize the correlations in benchmark scores to simplify the scaling law in terms of parameter count without harming prediction power by assuming that LLM performance is driven by a set of low-dimensional latent skills, such as reasoning and instruction following, which can be easily interpreted. Furthermore, we hypothesize that these latent skills are similarly influenced by computational resources, such as model size and training tokens, across different LLM families, with the key distinction being each family’s efficiency in converting compute into skill levels–which can be estimated with only one model per family. In summary, our main contributions are
@@
  • Providing a theoretical result regarding the identification of Sloth’s parameters and empirically demonstrating that our scaling laws can (i) accurately predict the performance of large models in 12 prominent LLM benchmarks and (ii) provide interpretable insights into LLM skills and scaling behavior.
  • Demonstrating how predicted latent skills can be used to predict (i) model performance in complex downstream tasks that involve coding and emotional intelligence and (ii) LLM behavior with scaled test-time compute.
@@ 1.1 Related work
Scaling laws for deep neural networks: In recent years, researchers have studied scaling laws from different angles. Rosenfeld et al. (2019) provides experimental scaling laws that predict model loss as a function of training set size, model width, and model depth. Likewise, Kaplan et al. (2020) establishes scaling laws that primarily measure loss (perplexity) and not accuracy on downstream tasks or benchmarks. Motivated by the presence of hard limits on the size of trainable data sets but a hypothetical unlimited ability to scale models, the authors of Muennighoff et al. (2023) establish scaling laws in constrained data settings. They find that perhaps unsurprisingly, increasing computing provides diminishing returns if data does not scale. Gadre et al. (2024) addresses the gap between the assumptions in scaling laws and how training is performed in practice; in particular, they construct scaling laws that both perform well in the over-training regime and predict performance on downstream tasks. In a similar but distinct direction, some works try not only to estimate scaling laws but also respond to the following strategic question: “Given a fixed FLOPs budget, how should one trade-off model size and the number of training tokens?” For example, Hoffmann et al. (2022) provides a partial answer, introducing the celebrated family of Chinchilla scaling laws and finding that training tokens and parameter size should roughly scale together. This contrasts with the older work of Kaplan et al. (2020) that provides a series of power laws that imply that simply increasing parameter count will provide good returns. Each of these referenced works trains models with a particular pretraining setting (e.g., architecture) at various sizes and ultimately seeks to predict test loss. Our focus is distinct, we fit scaling laws on existing benchmark data of multiple model families and predict LLM benchmark performance with minimal amount of data on the new family being predicted. Even though Allen-Zhu & Li (2024), study how LLMs can retain knowledge depending on their scale, the closest related works are Owen (2024); Ruan et al. (2024); Gadre et al. (2024); we will provide a detailed comparison with their work throughout the paper.
@@
LLMs latent skills: Given that the performance of large language models (LLMs) in different and diverse benchmarks is correlated, it makes sense to think that those models have some low-dimensional latent skills that are reflected in downstream tasks. In this direction, Ilić (2023) extracts a general intelligence factor (“g-factor”) for LLMs using the Open LLM Leaderboard (Beeching et al., 2023) and GLUE (Wang et al., 2018) using factor analysis. They also verify that this “g-factor” positively correlates with model size. In a similar direction Burnell et al. (2023) uses HELM (Liang et al., 2022) data to reveal that LLM intelligence may be constituted by three distinct, yet correlated factors. They also verify a positive correlation between model size and these latent skills, yet they do not propose a formal scaling law. In their study, the authors do not account for the training set size or model family information, leading to a poor fit of the regression model; this leaves good extrapolation as an open problem we address. In Kipnis et al. (2024), a unidimensional item response theory model is applied to each one of the 6 (filtered) benchmarks of the Open LLM Leaderboard. A factor analysis on the skill parameters shows that the main factor (carrying  80% of the data variability) is highly correlated with the “grand” (average) score of LLMs. In a related but different direction, Maia Polo et al. (2024a, b) show that inferring low-dimensional latent skills of LLMs can make model evaluation much more efficient, saving up to 140x in computing power. In this work, we explicitly model LLM skills as a function of computing resources, which enables the creation of accurate and interpretable scaling laws for benchmark performances.
@@ 3.1 Model architecture
We present a novel scaling law called Sloth, which introduces several modifications to (2.2). The key innovation of Sloth lies in its explicit modeling of the correlation structure between benchmarks, resulting in improved predictive accuracy and interpretability. Moreover, Sloth proposes that (i) LLM capabilities should scale with computing resources similarly across families up to an efficiency factor, (ii) benchmark performance can depend on $s$ and $t$ not only through the total number of FLOPs, and (iii) that the function $\sigma$ can also be learned in cases in which predictive performance is important. We detail these points.
Inspired by the latent skills (e.g., reasoning, language modeling, instruction following) inferred from benchmark data in Burnell et al. (2023); Ilić (2023); Ruan et al. (2024); Gor et al. ; Kipnis et al. (2024); Maia Polo et al. (2024a, b), we propose creating a scaling law for LLMs skills by leveraging the correlation structure of the benchmarks; for example, we model how the construct “reasoning” scales with compute instead of modeling benchmarks scores directly. The two major advantages of this approach are better performance prediction since we have fewer parameters to fit (reducing overfitting) and extra interpretability/insights. Concretely, we model $\eta_{ij}(s,t)$’s simultaneously for benchmarks $j\in\mathcal{J}$ as each being a linear combination of the same low-dimensional latent skills $\theta_{i}(s,t)\in{\mathbf{R}}^{d}$ plus a bias term $b\in{\mathbf{R}}^{J}$, where $d\ll J=|\mathcal{J}|$. Denote $\eta_{i}(s,t)\in{\mathbf{R}}^{J}$ as the vector of $\{\eta_{ij}(s,t)\}_{j\in\mathcal{J}}$. Mathematically, we have
@@
One can see that $\Lambda\in{\mathbf{R}}^{J\times d}$ encodes the correlation structure between the benchmarks; in particular, it tells us which benchmark measures overlapping (or distinct) skills. Interestingly, our model has a strong connection with factor analysis (FA) models, which we elaborate on in detail in Appendix D. In FA, the matrix $\Lambda$ is known as factor loadings while $\theta_{i}(s,t)$ are known as factors.
@@
Note that (i) the intercept parameter $\alpha_{ik}$ is indeed family-dependent while each skill slope is shared across families and (ii) $\theta_{i}$ can depend on $s$ and $t$ not only through $c(s,t)$. In economic terms, the intercept term $\alpha_{ik}$ can be interpreted as an efficiency measure of the family $i$ in converting compute to performance for skill $k$ and, in practice, will absorb all hidden factors specific to family $i$ such as data quality, post-training factors, etc.. We note that the interaction term in ($\log(s)\log(t)$) accounts for the fact that the impact of $\log(s)$ and $\log(t)$ on skills might depend on each other; in Appendix E, we show some evidence that this is indeed the case. Additionally to the changes in $\eta_{ij}$, we propose making the activation function $\sigma$ trainable and specific to each benchmark $j$ if needed. To that end, we adopt a semi-parametric single-index model using neural networks (Bietti et al., 2022). To make the results more behaved if (out-of-support) generalization is needed, we assume $\sigma_{j}:{\mathbf{R}}\to[0,1]$ is given by a monotonic (increasing) neural network, which can be achieved by constraining its weights to be non-negative (Sill, 1997) and its last activation function to be sigmoid. We note, however, that one can always forgo training of the link function and instead assume a sigmoid structure as this simplifies model fitting and may make the model more interpretable. We give more details about model fitting in Appendix C.
@@ 3.2 Interpretability and practical considerations post model fitting
Ideally, to make reasonable interpretations of models like Sloth, we need its parameters to be identifiable, i.e., the map connecting $\eta_{ij}(s,t)$’s and model parameters should be bijective. Unfortunately, as in all exploratory factor models, this is not the case. However, we theoretically show in Appendix B that the model parameters are identifiable up to some parameter transformations. This observation allows us to find a configuration that makes the model more interpretable by mirroring a standard approach used in factor analysis, e.g., in Chen et al. (2019)’s applications. First, we fit Sloth without any constraints on the covariance of skills obtaining the estimates $(\hat{\Lambda},\hat{b},\hat{B})$. Second, we find the matrix $A\in{\mathbf{R}}^{d\times d}$ such that the skills $X\hat{B}A$ have covariance identity, update $\hat{B}\leftarrow\hat{B}A$, and update $\hat{\Lambda}\leftarrow\hat{\Lambda}(A^{\top})^{-1}$ so the model outputs remains unchanged, because $\hat{\Lambda}(X\hat{B})^{\top}=\hat{\Lambda}(A^{\top})^{-1}(X\hat{B}A)^{\top}$. Third, we find a matrix $M\in{\mathbf{R}}^{d\times d}$ such that $\hat{\Lambda}M$ is easily interpretable (e.g., it is a sparse matrix); there are different methods to find $M$ and, in this paper, we use the Geomin (Yates, 1987; Chen et al., 2019) oblique rotation method to find a suitable $M$ using the Python package FactorAnalyzer (Biggs, 2019). We then update $\hat{\Lambda}\leftarrow\hat{\Lambda}M$ and, to make the model invariant, we also update $\hat{B}\leftarrow\hat{B}(M^{\top})^{-1}$; the resulting skills are still guaranteed to have unitary standard deviations, so their covariance equals their correlation. Finally, we standardize the columns of the skills $X\hat{B}$ to have zero mean, while keeping the correlation structure unchanged. This last step implies that $\hat{b}$ must be translated to make the model invariant.
@@ 4 Sloth in practice
In this section, we present experimental results that provide evidence of the usefulness of Sloth. We perform experiments on a set of twelve benchmarks and state-of-the-art LLM families, including LLaMa 3 (Dubey et al., 2024), Qwen 2 (Yang et al., 2024), and Yi 1.5 (Young et al., 2024). We explore the following applications: (i) benchmark performance prediction for larger models from a specific LLM family, (ii) interpretability of the scaling of skills (can help practitioners allocate resources based on the skills of interest), and (iii) downstream tasks performance prediction.
@@ 4.1 Data
We expand the dataset made available by Ruan et al. (2024), including more models from the HuggingFace Open LLM leaderboard v1 (Beeching et al., 2023) and v2 (Fourrier et al., 2024). In our extended dataset, we have a total of 30 families [^1] If we consider that instruct and base models are from different families, we end up with 53 families., which 28 are on v1 of the Open LLM Leaderboard and 17 families measured on v2 of the Open LLM Leaderboard. Furthermore, there are 15 families at the intersection of the two versions. Furthermore, we collect data and present results on the performance of a variety of instruction-tuned versions of the base models we consider. As far as we are aware, our dataset is the most comprehensive among prior works on benchmark data scaling laws. Please check Appendix G for details on the included models.